An Earth-shattering Work Significantly Advancing the State of the Art

by

An Author, BCompSc

Thesis

Submitted by An Author

for fulfillment of the Requirements for the Degree of

Doctor of Philosophy (0190)

Supervisor: Dr. Joe Bloggs Associate Supervisor: Dr. Fred Nurke

Clayton School of Information Technology

Monash University

May, 2006 \begin{matrix}c \\ ◯ \end{matrix} Copyright

by An Author I luv youse all

iii

Contents

List of Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v List of Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vi Abstract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . vii Acknowledgments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix 1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Literature Review . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Figures and Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.2.1 Referencing test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Appendix A This appendix should get a letter . . . . . . . . . . . . . . . . 9 Appendix B Simulation Source Code . . . . . . . . . . . . . . . . . . . . . . . 11 Vita . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Last Thing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

List of Tables

3.1 An example of a table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

List of Figures

3.1 An example of a figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

An Earth-shattering Work Significantly Advancing the State of the Art

An Author, BCompSc

                        authora@csse.monash.edu.au

Monash University, 2006

Supervisor: Dr. Joe Bloggs

                        bloggsj@csse.monash.edu.au

Associate Supervisor: Dr. Fred Nurke

                         nurkef@csse.monash.edu.au

Abstract

This thesis introduces … Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat.

An Earth-shattering Work Significantly Advancing the State of the Art

Declaration

I declare that this thesis is my own work and has not been submitted in any form for another degree or diploma at any university or other institute of tertiary education. Information derived from the published and unpublished work of others has been acknowledged in the text and a list of references is given.

An Author May 10, 2006

Acknowledgments

I would like to thank everyone who helped to make this possible. It has been an incredible journey of self-discovery, and I love every last one of you…

An Author

Monash University May 2006 x

Chapter 1

Introduction

Here I introduce the research area and my contributions to it.

Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat.

Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat. Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi.

Nam liber tempor cum soluta nobis eleifend option congue nihil imperdiet doming id quod mazim placerat facer possim assum. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat. Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis.

At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, At accusam aliquyam diam diam dolore dolores duo eirmod eos erat, et nonumy sed tempor et et invidunt justo labore Stet clita ea et gubergren, kasd magna no rebum. sanctus sea sed takimata ut vero voluptua.

2 CHAPTER 1. INTRODUCTION est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat.

Literature Review

In this chapter I will demonstrate some of the extended citation capabilities provided by the natbib package (Daly, 1999). It replaces the standard L^{A}TEX \cite command with two basic forms of citation command: \citep and \citet, as well as providing several other very useful ones.

The \citep command is best used when placing a citation at the end of sentence or phrase (as above). In the natbib documentation, this is referred to as a parenthetical citation.^{1}

When you want to refer to the authors of a particular work, typically at the start of a sentence, a parenthetical citation is not appropriate. This is particularly so if you are using a numerical or symbolic citation style. You should not start a sentence with

[2] says that this is most certainly …

In such situations you really need to give the authors’ names. The \citet command produces textual citations, which allows you to produce things like:

Ade (1983) describes a means by which textures may be characterized … another approach is given in De Bonet and Viola (1998).

Atick et al. (1996) note that humans have little or no difficulty in perceiving shape, yet find it extremely difficult to describe what they perceive.

Note that an abbreviated version of the authors’ names has been produced in the third example above. It is often desirable to have the full list of authors’ names given when a work is first cited, and an abbreviated list thereafter. This can be achieved by passing the longnamesfirst option to natbib when the package is used. This will produce an initial citation like:

Atick, Griffin and Redlich (1996) note that humans have little or no difficulty in perceiving shape, yet find it extremely difficult to describe what they perceive.

Both the \citep and \citet can take two optional arguments. If just one is provided, its text will appear as a “post-note” after the citation details. If two arguments are provided, the first defines a pre-note, and the second a post-note. Here is an example:

\citep[Ch.~3]{AaK1989} … (Aarts and Korst, 1989, Ch. 3)

\citep[see][Ch.~3]{AaK1989} … (see Aarts and Korst, 1989, Ch. 3)

1For ease of conversion from exisiting L^{A}TEX documents, you might find it useful to place \renewcommand{\cite}{\citep} in the preamble of the document, since any existing standard \cite commands should almost certainly be treated as \citep.

These examples only scratch the surface of what the natbib package can do. To discover the full power of the package, see the documentation at CTAN (Daly, 1999). You probably already have it on your system. Try locate natbib.dvi at the command line.

Chapter 3

Figures and Tables

Here we will test that references to figures and tables work correctly.

3.1 Figures

Figure 3.1: An example of a figure.

See Figure 3.1.

3.2 Tables

See Table 3.1.

3.2.1 Referencing test

See Table 3.1 and Figure 3.1.

23121 1212 232 cat frog dog Table 3.1: An example of a table 6 CHAPTER 3. FIGURES AND TABLES

Method

Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat.

Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat. Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis at vero eros et accumsan et iusto odio dignissim qui blandit praesent luptatum zzril delenit augue duis dolore te feugait nulla facilisi.

Nam liber tempor cum soluta nobis eleifend option congue nihil imperdiet doming id quod mazim placerat facer possim assum. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat. Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis.

At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, At accusam aliquyam diam diam dolore dolores duo eirmod eos erat, et nonumy sed tempor et et invidunt justo labore Stet clita ea et gubergren, kasd magna no rebum. sanctus sea sed takimata ut vero voluptua. est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat.

Consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Nam liber tempor cum soluta nobis eleifend option congue nihil imperdiet doming id quod mazim placerat facer possim assum. Lorem ipsum dolor sit amet, consectetuer adipiscing elit, sed diam nonummy nibh euismod tincidunt ut laoreet dolore magna aliquam erat volutpat. Ut wisi enim ad minim veniam, quis nostrud exerci tation ullamcorper suscipit lobortis nisl ut aliquip ex ea commodo consequat.

Duis autem vel eum iriure dolor in hendrerit in vulputate velit esse molestie consequat, vel illum dolore eu feugiat nulla facilisis.

At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, At accusam aliquyam diam diam dolore dolores duo eirmod eos erat, et nonumy sed tempor et et invidunt justo labore Stet clita ea et gubergren, kasd magna no rebum. sanctus sea sed takimata ut vero voluptua. est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat.

Consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet. Lorem ipsum dolor sit amet, consetetur sadipscing elitr, sed diam nonumy eirmod tempor invidunt ut labore et dolore magna aliquyam erat, sed diam voluptua. At vero eos et accusam et justo duo dolores et ea rebum. Stet clita kasd gubergren, no sea takimata sanctus est Lorem ipsum dolor sit amet.

Appendix A

This appendix should get a letter

An appendix before the backmatter gets an automatically generated letter by which it can be referred to. This is Appendix A.

10 APPENDIX A. THIS APPENDIX SHOULD GET A LETTER

Appendix B

Simulation Source Code

You may want to investigate the lgrind program and package if you wish to include source code in your thesis 12 APPENDIX B. SIMULATION SOURCE CODE

Vita

Publications arising from this thesis include: Author, A. and Bloggs, J. (2002), A really catchy title. In The 31st International

Conference on Non-specific Computing. Capital City, Country. Bloggs, J. and Author , A. (2002), A very much longer and significantly less catchy

title. in Workshop on A Research Area. Springfield, USA.

Permanent Address: Clayton School of Information Technology Monash University Australia

This thesis was typeset with L^{A}TEX 2ε^{1} by the author.

1L^{A}TEX 2ε is an extension of L^{A}TEX. L^{A}TEX is a collection of macros for TEX. TEX is a trademark of the American Mathematical Society. The macros used in formatting this thesis were written by Glenn Maughan and modified by Dean Thompson and David Squire of Monash University.

14 VITA

References

Aarts, E. and Korst, J. (1989). Simulated annealing and Boltzmann machines, John Wiley and Sons, New York.

Ade, F. (1983). Characterisation of textures by “eigenfilters”, Signal Processing 5: 451– 457.

Atick, J. J., Griffin, P. A. and Redlich, A. N. (1996). The vocabulary of shape: principal shapes for probing perception and neural response, Neural Computation 7(1): 1–5.

Daly, P. W. (1999). Natural sciences citations and references (author–year and numerical schemes), L^{A}TEX2e package documentation. (last accessed February 7, 2004). URL: http://www.ctan.org/tex-archive/macros/latex/contrib/natbib/

De Bonet, J. S. and Viola, P. (1998). Texture recognition using a non-parametric multiscale statistical model, Proceedings of the 1998 IEEE Conference on Computer Vision and Pattern Recognition (CVPR’98), Santa Barbara, California, USA. URL: http://www.ai.mit.edu/˜ jsd/research/publications/1998/DeBonet-CVPR98.pdf 16 REFERENCES

Last Thing

This sort of appendix has no letter.